OpenStudy (anonymous):

7r-7=2r+18 please show me how to solve it

OpenStudy (anonymous):

First you need to isolate your variable, this means get the r's on to one side so you can equal it to a number

OpenStudy (secret-ninja):

@magy Do you still need help?

OpenStudy (secret-ninja):

7r-7=2r+18 Move the "-7" to the other side of the equal sign. Change the minus sign to a plus sign. 7r=2r+18+7 18+7 is 25. So... 7r=2r+25

OpenStudy (anonymous):

yes thank you the answer is 5

OpenStudy (secret-ninja):

You then move the 2r to the other side. 7 r - 2r is 5r so now you have 5r=25 and you divide 5 by r and 25 and you get 5.

OpenStudy (secret-ninja):

Oh okay lol no problem. ^-^